# DREAM: Dynamic Red-teaming across Environments for AI Models

### v1

@@ 5.1.3 Our Attack Framework
The action space $\mathcal{A}$ for the Conductor is a comprehensive Atom Attack Library built upon 349 distinct digital environments from SafetyBenchmark [yin2025safeagentbenchbenchmarksafetask]. Our multi-agent generation pipeline produced 1,986 unique atom attacks, systematically distributed across three adversarial roles as summarized in Table V.
TABLE V: Detailed Composition of the Atom Attack Library, Categorized by Adversarial Roles and Coverage Across Digital Environments.
@@
Validation of the Action Space. To ensure the rigor and challenge of our evaluation, we conducted a deep statistical analysis of the generated attack library. As summarized in Table VI, our library is not merely voluminous but methodically constructed to probe specific safety deficits.
TABLE VI: Statistical Analysis of the Atom Attack Library: Prevalence of Top Risk Categories and Failure Modes.
Type Category / Mode Description Count Pct.(%)
@@
The library demonstrates high complexity, with over 52.6% of attacks involving multiple concurrent risk categories. In terms of risk impact, it heavily targets high-stakes vulnerabilities, with Sensitive Data Leakage (R1, 23.4%) and Property Loss (R2, 19.1%) being the most frequent categories.
Crucially, the distribution of failure modes reveals that the attacks are designed to challenge deep reasoning capabilities rather than simple compliance. The most prevalent failure mode, Ignoring Implicit Risks (M5, 29.2%), directly tests an agent’s ability to interpret unstated contextual dangers, a key component of preventing the ”Contextual Fragility” we investigate. Furthermore, 15.3% of attacks exploit Excessive Trust in Tool Results (M9). This empirically validates the efficacy of our Seeder agents, demonstrating their ability to manipulate the environment and mislead target agents into unsafe actions through deceptive information.
This extensive and diverse library allows the C-GPS algorithm to dynamically construct complex, causally-linked attack chains by leveraging the stateful knowledge maintained in the Cross-Environment Adversarial Knowledge Graph (CE-AKG).
5.2 Experiment Results
Overall Agent Performance
We present the comprehensive evaluation results in Table IV, showing the impact of our framework’s core innovations on current agent defenses. The effectiveness of our Contextualized Guided Policy Search (C-GPS) is clear in the Attack Success Rate column. By constructing long, causally-linked attack chains, C-GPS bypasses simple, single-turn safety measures, achieving success rates exceeding 70% for 8 out of the 12 models. This shows that modern agents struggle to track long-term malicious intent, a vulnerability central to the “domino effect” that C-GPS simulates.
The results also highlight a systemic failure in contextual reasoning, which our Cross-Environment Adversarial Knowledge Graph (CE-AKG) exploits. The low Contextual Isolation scores across most models—for instance, 37.82 for qwen3-235b-a22b and 44.90 for gemini-2.5-flash—quantify the “contextual fragility” that CE-AKG reveals. These agents struggle to prevent information and context from crossing environmental boundaries, letting our Conductor pivot and escalate attacks in ways not previously tested. Ultimately, the low Overall Defense Score for most agents results from the synergy between these two innovations, showing that agents are poorly equipped to handle adversaries that maintain state and pivot between contexts—a threat model that DREAM evaluates systematically for the first time.
5.2.1 Detailed Breakdown of Vulnerabilities by Model
To scrutinize the specific safety profiles of different agent architectures, we present a granular breakdown of Attack Success Rates (ASR) across all 12 evaluated models. Consistent with the overall rankings, we categorize models into proprietary and open-source groups. In these tables, lower ASR ($\downarrow$) indicates better defense, and the best performance in each category is highlighted in bold.
Risk Category Analysis. Table VII details the ASR across the eight risk categories. The data reveals a systemic vulnerability in Data Leakage (R1), which remains consistently high ($75\%$) for the majority of models. Notably, while claude-sonnet-4.5 achieves the best average defense (48.31%), grok-4 demonstrates superior resilience specifically in preventing data leakage (55.43%) and misinformative production (41.67%). In the open-source cohort, Vulnerable Code Generation (R7) is a critical weak point, with models like llama3.1-70b and gemini-nothinking reaching a 100% failure rate.
TABLE VII: Attack Success Rate (%) by Risk Category. Models are sorted by Overall Defense Score. Bold indicates the best performance (lowest ASR) in each column. $\downarrow$ denotes lower is better.
Model R1 R2 R3 R4 R5 R6 R7 R8 Avg.
Data ($\downarrow$) Prop. ($\downarrow$) Misinf ($\downarrow$) Harm ($\downarrow$) Legal ($\downarrow$) Avail ($\downarrow$) Code ($\downarrow$) Prod ($\downarrow$) ($\downarrow$)
@@ 5.2.1 Detailed Breakdown of Vulnerabilities by Model
Kimi-K2-0711-preview 83.33 70.59 54.24 69.86 69.23 60.58 80.00 57.89 70.02
llama3.1-70b 76.30 78.81 69.05 66.04 72.22 64.86 100.0 75.00 74.06
qwen2.5-72b 88.57 67.52 72.73 62.07 57.97 63.01 80.00 100.0 70.74
deepseek-v3.1-think-250821 83.91 67.14 79.41 67.54 73.54 66.13 83.33 72.09 72.86
Failure Mode Analysis. Table VIII examines cognitive failures. The data uncovers a universal operational vulnerability: Failure to Call Necessary Tools (M8). Almost all models exhibit extreme susceptibility here, with ASRs exceeding 90% for 9 out of 12 agents, indicating a tendency towards paralysis under attack. Furthermore, Ignoring Implicit Risks (M5) remains a prevalent deficit, with even advanced models like deepseek-v3.1 failing 72.4% of the time. In contrast, claude-sonnet-4.5 demonstrates exceptional adherence to constraints (M4, 37.0%) and parameter correctness (M6, 36.0%).
TABLE VIII: Attack Success Rate (%) by Failure Mode. M8 (Paralysis) is a universal weakness. Bold indicates the best performance (lowest ASR). $\downarrow$ denotes lower is better.
Model M1 M2 M3 M4 M5 M6 M7 M8 M9 Avg.
($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$) ($\downarrow$)
@@
Kimi-K2-0711-preview 40.0 100.0 100.0 69.7 68.7 60.9 77.8 93.0 73.2 70.0
llama3.1-70b 100.0 66.7 100.0 68.4 73.8 65.6 73.5 96.6 75.6 74.1
qwen2.5-72b 0.0 100.0 66.7 50.9 68.5 61.1 68.0 98.2 78.8 70.7
deepseek-v3.1-think-250821 66.7 87.5 100.0 68.6 72.4 63.4 69.7 95.0 73.9 72.9
@@ 5.2.2 The Domino Effect: Power of Long-Chain Attacks
Figure 3: “Domino Effect” in Attack Chain Length and Final Score Distribution. The mean score (solid line) shows steep, super-linear growth, significantly outperforming the exponential baseline (dashed line). This divergence is a direct result of the C-GPS algorithm’s ability to construct causally-linked attack sequences, demonstrating that the potency of attacks increases synergistically with chain length.
Figure 3 provides a powerful visualization of the “domino effect”, a direct consequence of our C-GPS algorithm’s strategic, long-chain planning. The analysis reveals a super-linear relationship between chain length and attack severity. The solid line, tracking the mean score, demonstrates a clear accelerating growth, significantly outpacing the simple exponential growth of the baseline (dashed line). This divergence is critical: it shows that C-GPS does not merely accumulate value from independent steps but creates synergistic attacks where each step amplifies the potential of the next. This trend is further supported by the score distributions (box plots), where the median score consistently rises and the interquartile range expands, indicating that longer chains not only have a higher average outcome but also unlock the potential for far more severe, high-scoring breaches. This empirically validates that the emergent vulnerabilities uncovered by DREAM are a product of stateful, causal chaining, a threat paradigm fundamentally invisible to single-turn evaluations.
Figure 4: The “Information Bridge” Effect in Attack Chain Length and Environment Count. The mean score (solid line) shows near-linear growth as attacks traverse more environments. This demonstrates the effectiveness of our CE-AKG in fusing disparate information to enable contextually rich, high-impact attacks. The widening distribution of scores (box plots) at higher environment counts indicates that cross-environment attacks unlock opportunities for more severe breaches.
@@ 5.2.3 The Information Bridge: Impact of Cross-Environment Attacks
A core innovation of our framework is the Cross-Environment Adversarial Knowledge Graph (CE-AKG), which acts as an “information bridge” to enable attack pivoting. Figure 4 provides clear empirical evidence of this mechanism’s power. We observe a strong positive correlation between the number of unique environments an attack chain traverses and the final score. The mean score (solid line) increases consistently up to four environments, underscoring that each cross-environment pivot provides the Conductor with novel attack vectors and enriched context, managed and provisioned by the CE-AKG.
This trend directly quantifies the risk of “contextual fragility”: agents fail to maintain a coherent safety posture when faced with inputs that leverage knowledge from external, seemingly unrelated contexts. The expanding interquartile range of the box plots at higher environment counts is particularly revealing. It signifies that while not all cross-environment attacks are catastrophic, this strategy is uniquely capable of unlocking the highest-severity vulnerabilities, which remain entirely hidden in single-context evaluations. This proves that the CE-AKG is fundamental to discovering systemic flaws in agent reasoning. The slight stabilization in mean score at five environments is a result of our experiment’s maximum chain length of five steps. Achieving five unique environments in a five-step chain requires every step to be a cross-environment pivot, which is not only rare but also often sub-optimal for developing an attack that requires multiple actions within a single compromised context. We anticipate this growth trend would continue with longer potential attack chains.
5.2.4 Statistical Validation of Synergistic Effects
For each chain length and environment count, we conducted a one-tailed Wilcoxon signed-rank test on the log-transformed differences between DREAM’s scores and the baseline scores for all 12 target agents. Our alternative hypothesis ($H_{a}$) stated that the median of these log-differences is significantly greater than zero. The results are presented in Table IX.
TABLE IX: Wilcoxon Signed-Rank Test Results vs. Exponential Baseline. The analysis compares DREAM’s final scores against the baseline ($1.19^{n}$) in log space. The sample size (LLM Nums) reflects the number of models that produced valid attack chains for the given configuration.
Configuration n LLM Nums Median Log-Diff Wilcoxon Stat. p-value
Chain Length 1 12 -0.1436 0.00 $1.0000$
2 12 0.8149 78.00 $0.001^{***}$
3 12 1.0460 78.00 $0.001^{***}$
4 12 1.2912 78.00 $0.001^{***}$
5 12 1.5625 78.00 $0.001^{***}$
Cross-Environment 1 12 0.1440 75.00 $0.0012^{**}$
2 12 1.0500 78.00 $0.001^{***}$
3 12 1.3144 78.00 $0.001^{***}$
4 12 1.4882 78.00 $0.001^{***}$
5 10 1.5032 55.00 $0.001^{***}$
$^{***}p<0.001$, $^{**}p<0.01$
The statistical analysis in Table IX provides a powerful confirmation of our core thesis. For all multi-step configurations (Chain Length $n\geq 2$ and Cross-Environment $n\geq 2$), the tests yield p-values of $p<0.001$. This allows us to decisively reject the null hypothesis, providing strong statistical evidence that the attack chains generated by DREAM are not merely cumulative but synergistically more effective.
@@ 5.2.5 Evaluation Stability and Reproducibility
The results, detailed in Table X, exhibit a high degree of stability. The variance across all five evaluation metrics for both models is remarkably low. For instance, the Overall Defense Score for gemini-2.5-flash remained tightly clustered around $41.06\pm 1.16$, while its Attack Success Rate was a consistent $76\pm 1\%$. This consistency demonstrates that despite the dynamic and stochastic nature of the attack generation process, the aggregated results produced by DREAM are robust and reproducible. This stability is crucial, confirming that our framework serves as a reliable benchmark for comparing the safety postures of different AI agents.
TABLE X: Evaluation Stability of the DREAM Framework Across Multiple Runs. We conducted three independent evaluation runs for two representative models, each comprising 100 generated attack chains initiated with different random seeds. The low variance in the results across all five metrics demonstrates the high stability and reproducibility of our benchmark.
Model Overall Defense Attack Success Damage Attack Contextual
Score $\uparrow$ Rate (%) $\downarrow$ Mitigation $\uparrow$ Observability $\uparrow$ Isolation (%) $\uparrow$
Model qwen3-235b-a22b
qwen3-235b-a22b (Run 1) 38.73 83 58.10 44.97 37.82
qwen3-235b-a22b (Run 2) 40.95 77 59.64 47.48 38
qwen3-235b-a22b (Run 3) 40.90 76 61.20 46.24 46.48
Model gemini-2.5-flash
gemini-2.5-flash (Run 1) 40.89 76 60.26 47.41 44.90
gemini-2.5-flash (Run 2) 42.29 75 61.68 49.09 50
gemini-2.5-flash (Run 3) 40.00 77 60.54 47.07 42
5.2.6 Impact of Conductor Agent Capability
@@
6 Discussion
6.1 On the Ineffectiveness of Static Defenses
To test common defense strategies against DREAM attacks, we ran an experiment by adding explicit safety instructions to a target agent’s system prompt. We chose gemini-2.5-flash as a representative mid-tier model and added instructions at the beginning to refuse harmful requests, such as those trying to access unauthorized data, bypass security controls, or cause damage.
The results were surprising: adding this static defense mechanism did not improve the agent’s security but actually made it worse. The model’s Overall Defense Score dropped from 40.89 to 36.62, while the Attack Success Rate increased from 76% to 83%. Nearly all defense metrics, including Damage Mitigation and Attack Observability, got worse.
This result shows a fundamental weakness of static defenses against dynamic, long-chain attacks. Our analysis suggests that C-GPS gradually builds a malicious context over multiple seemingly benign turns. This “contextual buildup” dilutes or overrides the initial safety instructions, making them ineffective by the time the critical exploit happens. The experiment confirms that agents have severe “contextual fragility” and struggle to track long-term malicious intent. While Contextual Isolation improved slightly (from 44.90 to 51.44), this was outweighed by declines in all other areas. Ultimately, this shows that defending against DREAM’s threats requires sophisticated, stateful security mechanisms that can reason about safety in evolving interactions, not just fixed rules.

### v2

@@ 4.1.3 Our Attack Framework
The action space $\mathcal{A}$ for the Conductor is a comprehensive Atom Attack Library built upon 349 distinct digital environments from SafetyBenchmark [43]. Our multi-agent generation pipeline produced 1,986 unique atom attacks, systematically distributed across three adversarial roles as summarized in Table 5.
To ensure the rigor and challenge of our evaluation, we conducted a deep statistical analysis of the generated attack library. As summarized in Table 6, our library is not merely voluminous but methodically constructed to probe specific safety deficits. This extensive and diverse library allows the C-GPS algorithm to dynamically construct complex, causally-linked attack chains by leveraging the stateful knowledge maintained in the Cross-Environment Adversarial Knowledge Graph (CE-AKG).
Finally, to rigorously assess both the lower and upper bounds of model resilience using this library, we employed a stratified experimental design: attack targets were evenly distributed (20% each) to target chain lengths from 1 to 5 steps. Crucially, a dynamic fallback mechanism was implemented: if a target length (e.g., 5 steps) failed, the attack collapsed to its longest successful sub-chain. This ensures our data reflects the maximum viable exploitation depth for each scenario.
Table 5: Atom attack library composition. Attacks categorized by adversarial roles and digital environment coverage.
@@
Table 6: Atom attack library statistics. Distribution of attacks across risk categories and failure modes.
Type Category / Mode Description Count Percentage(%)
@@
Table 7: ASR (%) across risk categories. Bold denotes best defense (lowest ASR, $\downarrow$) per category. Data Leakage (R1) remains a persistent systemic vulnerability across most architectures, while open-source models show catastrophic susceptibility to Vulnerable Code Generation (R7), often reaching 100% failure rates.
Model R1 $\downarrow$ R2 $\downarrow$ R3 $\downarrow$ R4 $\downarrow$ R5 $\downarrow$ R6 $\downarrow$ R7 $\downarrow$ R8 $\downarrow$ Avg. $\downarrow$
@@
Kimi-K2-Preview 83.33 70.59 54.24 69.86 69.23 60.58 80.00 57.89 70.02
Llama-3.1-70B 76.30 78.81 69.05 66.04 72.22 64.86 100.00 75.00 74.06
Qwen-2.5-72B 88.57 67.52 72.73 62.07 57.97 63.01 80.00 100.00 70.74
Deepseek-V3.1 83.91 67.14 79.41 67.54 73.54 66.13 83.33 72.09 72.86
Table 8: ASR (%) across failure modes. All agents show vulnerability to M8 (Paralysis). Bold denotes best performance ($\downarrow$). While “Tool Paralysis” (M8) constitutes a universal failure point with >90% ASR for most agents, Claude-Sonnet-4.5 demonstrates exceptional resilience in operational constraint adherence (M4) and parameter validation (M6).
[…]
@@
Kimi-K2-Preview 40.0 100.0 100.0 69.7 68.7 60.9 77.8 93.0 73.2 70.0
Llama-3.1-70B 100.0 66.7 100.0 68.4 73.8 65.6 73.5 96.6 75.6 74.1
Qwen-2.5-72B 0.0 100.0 66.7 50.9 68.5 61.1 68.0 98.2 78.8 70.7
DeepSeek-V3.1 66.7 87.5 100.0 68.6 72.4 63.4 69.7 95.0 73.9 72.9
4.2 Experiment Results
Overall Agent Performance. We present the comprehensive evaluation results in Table 4, showing the impact of our framework’s core innovations on current agent defenses. The effectiveness of our Contextualized Guided Policy Search (C-GPS) is clear in the Attack Success Rate column. By constructing long, causally-linked attack chains, C-GPS bypasses simple, single-turn safety measures, achieving success rates exceeding 70% for 8 out of the 12 models. This shows that modern agents struggle to track long-term malicious intent, a vulnerability central to the “domino effect” that C-GPS simulates.
The results also highlight a systemic failure in contextual reasoning, which our Cross-Environment Adversarial Knowledge Graph (CE-AKG) exploits. The low Contextual Isolation scores across most models, exemplified by 37.82 for Qwen3-235B-A22B and 44.90 for Gemini-2.5-Flash, quantify the “contextual fragility” that CE-AKG reveals. These agents struggle to prevent information and context from crossing environmental boundaries, thereby enabling our Conductor to pivot and escalate attacks in ways not previously tested. Ultimately, the low Overall Defense Score for most agents results from the synergy between these two innovations, demonstrating that agents are poorly equipped to handle adversaries that maintain state and pivot between contexts, a threat model that DREAM evaluates systematically for the first time.
4.2.1 Detailed Breakdown of Vulnerabilities by Model
Figure 3: Analysis of the “Domino Effect”. (a) Heatmap of average severity scores for successfully implemented attacks across 12 models. The x-axis represents the attack chain step count (1-5), and the color gradient transitions from light blue (low severity) to dark red (high severity). (b) Aggregate scores compared to an exponential baseline ($1.19^{x}$, dashed). The steep, super-linear growth and expanding quartiles visually confirm the synergistic amplification of attack potency as chains extend.
To evaluate the specific safety profiles of different agent architectures, we present a detailed breakdown of Attack Success Rates (ASR) across all 12 evaluated models. Following the categorization estabilshed in our overall rankings, we organize the analysis into proprietary and open-source groups. Throughout these tables, lower ASR indicates stronger defense capabilities, with the best-performing model in each category highlighted in bold.
Risk Category Analysis. Table 7 reveals striking disparities in vulnerability patterns across the eight risk categories. More alarmingly, Data Leakage (R1) emerges as a systemic weakness that transcends model architecture boundaries. The vulnerability landscape here is severe: over three-quarters of models exhibit ASRs exceeding 75%, suggesting fundamental challenges in establishing proper data boundaries during agentic operations.
Within this challenging context, two distinct defensive strategies emerge among proprietary models. While Claude-Sonnet-4.5 achieves the strongest overall protection with an average ASR of 48.31% across all categories, Grok-4 demonstrates a specialized defensive profile. Specifically, Grok-4 outperforms all competitors in two critical areas: it reduces data leakage attacks to a 55.43% success rate, which is nearly 20 percentage points below the category average, and limits misinformation production to 41.67%. This divergence suggests that architectural choices create trade-offs between general defense and specialized resistance.
The open-source ecosystem faces a qualitatively different threat landscape. Vulnerable Code Generation (R7) represents a catastrophic failure point, with models such as Llama3.1-70B and Gemini-NoThinking reaching complete breakdown: every single adversarial prompt successfully elicits harmful code, yielding 100% ASR. This perfect failure rate stands in stark contrast to the proprietary models’ performance in the same category, where even the weakest defender maintains ASR below 85%, highlighting a substantial capability gap in code security reasoning.
Failure Mode Analysis. Table 8 shifts focus from what agents are attacked on to how they fail cognitively. The most prevalent weakness identified is Failure to Call Necessary Tools (M8), which reveals a fundamental operational fragility. Nine out of twelve models demonstrate extreme susceptibility, with ASRs surpassing 90%. This widespread vulnerability indicates that when adversarial inputs introduce ambiguity or misdirection, the majority of agents default to cognitive paralysis, where they neither refuse the request nor activate the appropriate tool chain, effectively rendering their agentic capabilities inert.
Beyond tool-calling paralysis, Ignoring Implicit Risks (M5) constitutes a secondary but still critical vulnerability. Even sophisticated models struggle with contextual threat assessment: Deepseek-V3.1, despite its advanced reasoning architecture, fails to recognize implicit dangers in 72.4% of adversarial scenarios. This suggests that current safety training focuses disproportionately on explicit harm indicators while leaving agents vulnerable to subtly constructed threats. Against this backdrop of widespread failures, Claude-Sonnet-4.5 demonstrates exceptional robustness in constraint adherence and parameter validation. Its ASR for Constraint Violation (M4) stands at 37.0%, representing the lowest failure rate in this category and nearly halving the median performance. Similarly, for Incorrect Tool Parameter Usage (M6) , it achieves 36.0% ASR, indicating robust validation mechanisms that verify both the semantic appropriateness and syntactic correctness of tool invocations before execution. These results suggest that effective defense requires multi-layered validation extending beyond simple input filtering to encompass operational constraint checking and parameter sanity verification.
Figure 4: Analysis of the “Information Bridge” Effect. (a) Heatmap of severity scores vs. unique environment count. Empty cells indicate instances where the Conductor autonomously optimized for fewer pivots within the 5-step limit, rather than forcing maximum environment traversal. (b) Aggregate score distribution. The continuous upward trend against the baseline confirms that fusing context across diverse environments significantly amplifies attack severity.
@@ 4.2.2 The Domino Effect: Power of Long-Chain Attacks
Building upon the multi-stage experimental design and dynamic fallback mechanism detailed in Section 4.1.3, Figure 3 visualizes the resulting “Domino Effect.” By isolating the maximum viable exploitation depth for each scenario, we rigorously quantify how attack severity escalates with chain length. The heatmap in Figure 3(a) reveals a systematic vulnerability progression. Across all models, severity scores shift from “low risk” (blue, $\approx 1.5$) at Step 1 to “high danger” (red, >10>10) at Step 5. This progression is quantified in Figure 3(b), where the aggregate mean score (solid purple line) exhibits super-linear growth, significantly outpacing the exponential baseline ($1.19^{x}$). The widening variability at deeper steps indicates that long-chain interactions unlock high-impact vulnerabilities that are structurally inaccessible to single-turn attacks.
Synthesizing these results with Table 4 reveals an interesting insight: stronger general defense does not guarantee lower severity upon breach. For instance, Claude-Sonnet-4.5 achieves the highest Overall Defense Score (67.30) in Table 4, reflecting its robustness against most attempts. However, Figure 3(a) shows that when a long-chain attack does succeed against Claude (Step 5), it yields the highest severity score (14.4) among all models. This suggests a “Capability-Risk Paradox”: while advanced models are harder to compromise, their superior reasoning capabilities, once exploited via the Domino Effect, can be leveraged to cause significantly greater damage than less sophisticated models.
@@ 4.2.3 The Information Bridge: Impact of Cross-Environment Attacks
By synthesizing context from diverse sources via the CE-AKG, the Conductor executes attacks across multiple domains. Figure 4(a) illustrates this impact: the intensification of severity hot-spots at higher environment counts signals a distinct escalation in risk. Although data at 5 environments is limited due to the Conductor’s autonomous preference for efficiency over forced exploration, the aggregate trend in Figure 4(b) reveals a clear upward trajectory linking environmental diversity to attack scores. The increasing mean score quantifies Contextual Fragility, demonstrating that agents fail to maintain safety boundaries when inputs combine information from different contexts. Furthermore, the widening score distribution indicates that cross-environment pivots raise the upper bound of attack severity, unlocking critical vulnerabilities that are inaccessible to single-environment assessments.
Table 9: Stability analysis across three evaluation runs. Low variance demonstrates consistent DREAM performance.
Model / Run Overall Defense Attack Success Damage Attack Contextual
[…]
Qwen3-235B-A22B
     Run 1 38.73 83.00 58.10 44.97 37.82
     Run 2 40.95 77.00 59.64 47.48 38.00
     Run 3 40.90 76.00 61.20 46.24 46.48
Gemini-2.5-Flash
     Run 1 40.89 76.00 60.26 47.41 44.90
     Run 2 42.29 75.00 61.68 49.09 50.00
     Run 3 40.00 77.00 60.54 47.07 42.00
4.2.4 Impact of Conductor Agent Capability
@@ 4.2.5 Evaluation Stability and Reproducibility
The results, detailed in Table 9, exhibit a high degree of stability. The variance across all five evaluation metrics for both models is remarkably low. For instance, the Overall Defense Score for Gemini-2.5-Flash remained tightly clustered around $41.06\pm 1.16$, while its Attack Success Rate was a consistent $76\pm 1\%$. This consistency demonstrates that despite the dynamic and stochastic nature of the attack generation process, the aggregated results produced by DREAM are robust and reproducible. This stability is crucial, confirming that our framework serves as a reliable benchmark for comparing the safety postures of different AI agents.